OpenStudy (anonymous):

Solve the following system of equations using the substitution method. Enter your answers as (x,y) pairs, separated by commas. If the system is inconsistent, type INCONSISTENT. If the system is dependent, type DEPENDENT. x+3y=10 xy=7

OpenStudy (anonymous):

x + 3y = 10 xy = 7 --> x = 7/y sub 7/y in for x in the first equation 7/y + 3y = 10 (multiply by 1/7) (1/7)7/y + (1/7)3y = (1/7)10 y + 3/7y = 10/7 7/7y + 3/7y = 10/7 10/7y = 10/7 y = (10/7) / (10/7) y = 1 now sub 1 in for y x + 3y = 10 x + 3(1) = 10 x + 3 = 10 x = 10 - 3 x = 7 check... xy = 7 (1)(7) = 7 7 = 7 (correct) check... x + 3y = 10 7 + 3(1) = 10 7 + 3 = 10 10 = 10 (correct) ANSWER : (7,1) Dependent